Solve 5y – 8 + y = 9y + 23 – 6y
Answer by unlockmath(1688)

Hello,
The whole trick to Algebra is getting the variable (y) on one side of the equal sign. With this one 5y – 8 + y = 9y + 23 – 6y
First combine like terms to get:
6y-8=3y+23 Now subtract 3y and add 8 to both sides to get:
3y=31 Divide both sides by 3 to get:
y=31/3
Plug this in to the original and it will work out.
Make sense?
RJ
www.math-unlock.com